OpenStudy (princezombie):

An object is moving from point A to point B. At point A, the time t = 0 s, velocity v = 0 m/s, total force acting on the object F = 6000 N. by the time it reaches point B, velocity v = 4.95 m/s, total force F = 0 N. the equation for F is F = T - D, where T = const = 6000N and D = y * v^2, where y = 245 kg/m. Mass of the object is 300 kg. My problems are, how long does it takes for the object to move from point A to point B and what is the distance covered by the object (time, t at point B and distance from point A to B)?

OpenStudy (anonymous):

y= 244.8729721 (to get D = T). Get total F (integral of its equation) \[F _{total} = 6000 v - 81.6667 v ^{3}.(from 0 \to 4.95)\] \[F _{total} = 19794.86437\] \[dt = \frac{ M dv }{ F _{total} }\] dt = very small and this make sense from kinetic energy equation you can get work \[0.5Mv _{2} ^{2} - 0.5 M v _{2} ^{1} = W\] thus W = 7350.75 so dx = W/F = 0.371346318 m. and it is done :D

OpenStudy (anonymous):

i use y = 245 but i think it is wrong